OpenStudy (anonymous):

I WILL MEDAL AND FAN!!!!! Mr. Stewart is buying books and star charts for his astronomy group. His budget allows him to spend no more than $150. He plans to buy 8 books at $11 each. How many star charts could he also buy at $10 each? IS THE ANSWER 6?

OpenStudy (anonymous):

ye its 6 bc 8*11=88+6*10=148

OpenStudy (anonymous):

tnx for the medal i didnt deserve it this was pretty simple